OpenStudy (rainbow_rocks03):

HELP ME SOMEONE PLZ! Carol and Nora left the library at the same time. Carol biked east at a speed of 8 miles per hour. Nora biked west at a speed of 10 miles per hour. How far apart were Carol and Nora after 3 hours?

OpenStudy (rainbow_rocks03):



OpenStudy (rainbow_rocks03):

@Ilovecake

OpenStudy (anonymous):

tims 8 by 3 then 10 by 3 and then add your answers

OpenStudy (rainbow_rocks03):

8*3=24 and 10*3=30 30+24=54

OpenStudy (anonymous):

Yep so the answer is 54 miles

OpenStudy (ilovecake):

correct ^

OpenStudy (rainbow_rocks03):

thx for the help can you help me with 2 more?

OpenStudy (anonymous):

ya

OpenStudy (rainbow_rocks03):

let me open new ones

OpenStudy (anonymous):

K